Suppose that an economy is composed of three classes. Laborers, traders, and capital owners. The size of each of the classes is 60%, 30%, and 10% respectively. Laborers produce all the output in this economy by using capital borrowed from the capitalists and then using the services of traders to market their produce. For this, the laborers have to pay 40% and 20% of their produce to the capitalists and traders respectively.


a.Draw the Lorenz curve for the income distribution of this economy. Calculate the Gini Coefficient.


b.Suppose that the capitalist is taxed 10% of their income and this is distributed as transfers to the laborers. Draw the new Lorenz curve and compare income inequality between the market income distribution and the disposable income distribution.


1
Expert's answer
2021-10-25T08:46:52-0400

Solution:

a.). The Lorenz curve is a below:




 

Gini Coefficient = "\frac{A}{(A +B)}"


= "1 - (2\times 0.2) \times (0.133 + 0.267 + 0.20 + 0.40) = 0.6 \times1 = 0.6"


b.). b.). The new Lorenz curve is as below:




The income inequality between the market income distribution and disposable income distribution will slightly improve towards income equality.